OpenStudy (anonymous):

Find all values of θ in the interval 0° ≤ θ < 360° that satisfy the equation sin 2θ = sin θ. i know the answer just want to know how i do it thanks

OpenStudy (anonymous):

you need to rewrite \[\sin(2x)=2sin(x)cos(x)\]

OpenStudy (anonymous):

sin 2A = 2sinA cosA so 2sinAcosA=sinA 2cosA=1 cos A =1/2 fins all values of A that satisfy cos A =1/2

OpenStudy (asadkarim7):

i think its sin ^2 theta

OpenStudy (anonymous):

then go from there \[2sin(x)cos(x)-sin(x)=0\] \[sin(x)(2cos(x)-1)=0\] \[sin(x)=0\] or \[cos(x)=\frac{1}{2}\]

OpenStudy (anonymous):

first one gives x = 0 second one give x = 60 or x = 300

OpenStudy (anonymous):

since you are working in degrees

OpenStudy (anonymous):

2 sinθ cosθ n - sinθ = 0 sinθ (2 cosθ - 1) = 0 θ = 0, 180, 60, 300

OpenStudy (anonymous):

jimmyrep is right. i forgot that 180 works for sin(x)=0 as well. he is correct

OpenStudy (anonymous):

I still dont get it but thanks for trying i probably just need it spoken step by step ill ask tomorrow :]